1 A= Find A1017 and Justify your answer 1 ✓2 B = O 01 3 Find B” and Justify your answer. 0 3
Answer #1

Hey there stay home and stay safe

\small A=\begin{bmatrix} \frac{1}{\sqrt2} & -\frac{1}{\sqrt2}\\ \frac{1}{\sqrt2} & \frac{1}{\sqrt2}\\ \end{bmatrix}\\ A^2=\begin{bmatrix} \frac{1}{\sqrt2} & -\frac{1}{\sqrt2}\\ \frac{1}{\sqrt2} & \frac{1}{\sqrt2}\\ \end{bmatrix}\begin{bmatrix} \frac{1}{\sqrt2} & -\frac{1}{\sqrt2}\\ \frac{1}{\sqrt2} & \frac{1}{\sqrt2}\\ \end{bmatrix}=\begin{bmatrix} \frac{1}{2}-\frac{1}{2} & -\frac{1}{2}-\frac{1}{2}\\ \frac{1}{2}+\frac{1}{2} & -\frac{1}{2}+\frac{1}{2} \end{bmatrix}=\begin{bmatrix} 0 & -1\\ 1 & 0 \end{bmatrix}\\ A^4=A^2.A^2=\begin{bmatrix} 0 & -1\\ 1 & 0 \end{bmatrix}\begin{bmatrix} 0 & -1\\ 1 & 0 \end{bmatrix}\\ =\begin{bmatrix} 0-1 & 0+0\\ 0+0 & -1+0 \end{bmatrix}=\begin{bmatrix} -1 & 0\\ 0 & -1 \end{bmatrix}=(-1)\begin{bmatrix} 1 & 0\\ 0 & 1 \end{bmatrix}=(-1)I_2\\ A^{1017}=A^{1016}A=(A^4)^{254}A=(-1)^{254}I_2^{254}A=(1)I_2A=A=\begin{bmatrix} \frac{1}{\sqrt2} & -\frac{1}{\sqrt2}\\ \frac{1}{\sqrt2} & \frac{1}{\sqrt2}\\ \end{bmatrix}

....

\small B=\begin{bmatrix} \frac{1}{3} & \frac{1}{3}\\ 0 & \frac{1}{3} \end{bmatrix}=\frac{1}{3}\begin{bmatrix} 1 & 1\\ 0 & 1 \end{bmatrix}\\ B^2=\begin{bmatrix} \frac{1}{3} & \frac{1}{3}\\ 0 & \frac{1}{3} \end{bmatrix}\begin{bmatrix} \frac{1}{3} & \frac{1}{3}\\ 0 & \frac{1}{3} \end{bmatrix}=\begin{bmatrix} \frac{1}{3^2} & \frac{2}{3^2}\\ 0 & \frac{1}{3^2} \end{bmatrix}=\frac{1}{3^2}\begin{bmatrix} 1 & 2\\ 0 & 1 \end{bmatrix}\\ B^3=\frac{1}{3}\begin{bmatrix} 1 & 1\\ 0 & 1 \end{bmatrix}\frac{1}{3^2}\begin{bmatrix} 1 & 2\\ 0 & 1 \end{bmatrix}=\frac{1}{3^3}\begin{bmatrix} 1 & 1\\ 0 & 1 \end{bmatrix}\begin{bmatrix} 1 & 2\\ 0 & 1 \end{bmatrix}=\frac{1}{3^3}\begin{bmatrix} 1 & 3\\ 0 & 1 \end{bmatrix}\\ $So we claim that $\\ B^n=\frac{1}{3^n}\begin{bmatrix} 1 & n\\ 0 & 1 \end{bmatrix}\\ $To prove that we use principal of induction\\ Let it be true for n=k$\\ B^k=\frac{1}{3^k}\begin{bmatrix} 1 & k\\ 0 & 1 \end{bmatrix}\\

\small B^{k+1}=B^kB=B^n=\frac{1}{3^k}\begin{bmatrix} 1 & k\\ 0 & 1 \end{bmatrix}\frac{1}{3}\begin{bmatrix} 1 & 1\\ 0 & 1 \end{bmatrix}=\frac{1}{3^{k+1}}\begin{bmatrix} 1 & k+1\\ 0 & 1 \end{bmatrix}\\ $Then the statement is true for n=k+1 if it be true for n=k\\ Therefore by principal of induction$\\ B^n=\frac{1}{3^n}\begin{bmatrix} 1 & n\\ 0 & 1 \end{bmatrix}

And please leave a like. Thank you.